OpenStudy (anonymous):

Solve by Subsitution y=2x-1 3x+6y=4 Solve by Subsitution y=2x-1 3x+6y=4 @Mathematics

OpenStudy (anonymous):

3x + 6(2x-1)= 4 3x + 12x -6 = 4 15x=10 x=2/3

jimthompson5910 (jim_thompson5910):

3x+6y=4 3x+6(2x-1)=4 3x+12x-6=4 15x-6=4 15x=4+6 15x=10 x=10/15 x=2/3 From here, use this value of x to find y y = 2x-1 y = 2(2/3)-1 y = 4/3-1 y = 1/3

OpenStudy (anonymous):

Thanks!(: